OpenStudy (anonymous):

(Will give medal) If f(x) = log3 (x + 1), what is f−1(2)? Answer Selection A. 1 B. 8 C. 10 D. 27 I will point this out again. I am not asking to do a step by step with you on the comments section to learn to do this. I am on a time crunch, I am looking to get the answer and some of the work. I'm not trying to cheat and just get the answer because I know I will need to learn/master this lesson later. Thanks!

OpenStudy (anonymous):

Write it as: y = log3(x+1) switch x and y: x=log3(y+1) solve for y. This is your inverse function. Once you have that, plug in 2 for x, and calculate the answer. Small hint to make it easier: x = log3(y) solving for y gives: 3^x=y

OpenStudy (tkhunny):

Have you considered solving: \(2 = \log_{3}(x+1)\)

OpenStudy (anonymous):

Thanks for the help but guys please just stop dancing around the answer and answer it already! Please!

OpenStudy (anonymous):

I won't do your homework for you, but from what I said, you literally only have 3 steps to find your answer.

OpenStudy (tkhunny):

No one is dancing. You are to be learning. You have enough information. Use it. I already know how to do it. I won't learn anything by doing it.

OpenStudy (anonymous):

It's B.

OpenStudy (anonymous):

Thanks a bunch!